200 ml of h202 having volume strength 22.7 is mixed with 800 ml h2o2 solution having volume strength 45.4 are mixed together. the density of final h2o2 solution is 1.0224 g/ml
find 
1. final volume strength 
2. molality of final solution
3. w/v % of final solution

Grade:10

2 Answers

Nikhil
4 years ago
H2o2 ------  H20 + ½ O2 
volume strength of hydrogen peroxide is defined as amount of oxygen that can be produced by 1 litre of it...
Acc. New sign conventions 1mole gas = 22.7 ltr at STP
So Molarity of first sample is 2M and that of 2nd is 4M
so moles of h2o2 in 1st sample is 0.4 and in 2nd it is 4*800/1000 =3.2
after mixing voluume of final soln is1 ltr and total moles =3.6 this is also molarity 
so volume str = (3.6/2)*22.7 = 40.86 ltr (1.)
w/v% = weight of solute(h2o2)/volume of solution *100
         =3.6*34(molarmass of h2o2)/1000  *100
          =12.24% (3.)
molality = moles of solute/weight of solvent in kg
Now weight  of soln = density *volume = 1.0224*1000 =1022.4 gm
weight of h202 = 3.6*34= 122.4 gm
so weight of solvent = 1022.4-122.4 =900gm
             = 3.6/ 0.9 =4 (2.)
 
 
 
 
 
 
SAMBHAV MISHRA
4 years ago
@nikhil thank you very much bro.. god bless you..................................u r doig a nice job 
keep it up...